OpenStudy (anonymous):

You received a $100 gift certificate to a clothing store. The store sells t-shirts for $15 and dress shirts for $23. You want to spend no more than the amount of the gift certificate. You want to leave at most $5 of the gift certificate unspent. You need at least one dress shirt. What are all of the possible combinations of t-shirts and dress shirts you could buy? a. 1 t-shirt and 4 dress shirts b. 5 t-shirts and 1 dress shirt c. 3 t-shirts and 2 dress shirts d. 2 t-shirts and 3 dress shirts please help me solve this so i can pass my junior year.

OpenStudy (aripotta):

set up an equation, first

OpenStudy (anonymous):

how would i do that? gosh that makes me sound stupid...

OpenStudy (aripotta):

brb

OpenStudy (anonymous):

okay

OpenStudy (aripotta):

oh congrats lol

OpenStudy (anonymous):

AHAHAH i love the word it puts when you cuss

OpenStudy (aripotta):

yea, i know

OpenStudy (anonymous):

thats awesome

OpenStudy (anonymous):

hahah it missed one

OpenStudy (anonymous):

ahahahaahahah!

OpenStudy (anonymous):

i love this!! well im done bye bye

OpenStudy (opcode):

You received a $100 gift certificate to a clothing store. The store sells t-shirts for $15 and dress shirts for $23. You want to spend no more than the amount of the gift certificate. You want to leave at most $5 of the gift certificate unspent. You need at least one dress shirt. What are all of the possible combinations of t-shirts and dress shirts you could buy? a. 1 t-shirt and 4 dress shirts b. 5 t-shirts and 1 dress shirt c. 3 t-shirts and 2 dress shirts d. 2 t-shirts and 3 dress shirts $100 (Total Money)| $15 (per t-shirt)| $23 (per dress shirt)| $5 (Needs to be left.)| Okay so: 15 * 2 = 30 15 * 3 = 45 15 * 4 = 60 15 * 5 = 75 15 * 6 = 90 (Rates of t-shirts) 23 * 1 = 23 23 * 2 = 46 23 * 3 = 69 23 * 4 = 92 23 * 5 = 115 23 * 6 = 138 (Rates of dress-shirts) 1 t-shirt and 4 dress shirts (15) + (92) = 107 (Not quite it) 5 t-shirts and 1 dress shirt (75) + 23 = 98 ( Not quite it) 3 t-shirts and 2 dress shirts (45) + (46) = 91 (Not quite it) 2 t-shirts and 3 dress shirts (30) + (69) = 99 (Not quite it) (You want to leave at most $5 of the gift certificate unspent) I'd go with "3 t-shirts and 2 dress shirts (45) + (46) = 91 (Not quite it)."

OpenStudy (opcode):

@rachy Do you understand?

OpenStudy (anonymous):

yeah man i got it(:

OpenStudy (opcode):

Good glad to hear ^_^.

OpenStudy (anonymous):

thank ya

OpenStudy (opcode):

No problem :)

OpenStudy (anonymous):

whats your name?

OpenStudy (opcode):

I'm Opcode :)

OpenStudy (anonymous):

no your real name

OpenStudy (anonymous):

im rachel... kinda obviously...

OpenStudy (opcode):

Oh. Then I'm Kami.

OpenStudy (anonymous):

really where do you live?

OpenStudy (opcode):

I live in America. Why?

OpenStudy (anonymous):

idk how is that your name? thats kinda weird... no offence

OpenStudy (opcode):

It's a nickname.

OpenStudy (anonymous):

oh. whats your real name?

OpenStudy (opcode):

I'm not keen on giving away my name. Sorry.

OpenStudy (anonymous):

sorry... i guess its just your first name i dont see what the big deal is but okay thats cool i guess. i said i guess twice thats weird but whatevs.

OpenStudy (anonymous):

and "not too keen" what are you like 40? thats weird. idk anyone who says that and im 17